. The length of time X that the battery in Stella’s phone holds enough charge to operate is normally distributed with a mean μ = 25.6 hours and a standard deviation of σ = 0.32 hours. Stella forgot to charge her phone yesterday, and it’s been now 26 hours since the phone was charged. What’s the probability that the phone will operate?

Hint. The phone will operated if this particular time the battery lasted 26 or more hours


1
Expert's answer
2021-05-07T11:42:58-0400

"P(X>26)=P(Z>\frac{26-25.6}{0.32})=P(Z>1.25)=1-P(Z<1.25)=0.1056."